# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v1

@@ 3.5 Simulated or Hybrid Settings
It may not be always the case that the model will always observe a particular kind of incremental scenario, for example, domain incremental condition. The CL model developed for handling domain incremental conditions may not perform well for a class incremental condition. Thus, there are attempts to evaluate the same model on different CL scenarios; however, a separate evaluation is followed for each kind of incremental CL scenario. In contrast, having a set of episodes with a mix of CL scenarios is close to real-life conditions. Thus it becomes important to also test the performance in hybrid settings. An attempt in direction is made by Sadafi et al. (2023) where other than pure domain incremental and class incremental, a hybrid incremental scenario is also considered. The authors create a sequence of episodes with class and domain-incremental cases on multi-site WBC classification datasets. They create a long sequence of episodes with a total of 12 episodes. First, they set up a class incremental scenario within the Matek-19 dataset by adding 3-4 classes in each subsequent episode up to a total of four episodes. After this INT-20 dataset was considered in the same fashion, contributing 4 more episodes. After this, a third domain i.e., a new dataset Acevedo-20 was introduced in 4 episodes in class incremental fashion.
@@
Further, there are some attempts to curate shifts in datasets and various incremental scenarios instead of collecting actual datasets reflecting the situation. For example, Kaustaban et al. (2022) simulated domain shift in H&E stain due to distinct appearance caused by different staining protocols. This dataset is designed to simulate real-world scenarios, where data shifts occur due to differences in scanners, stainers, reagents, and other factors. They considered 9 classes from CRC Kather et al. (2019) which is a H&E stain-based colorectal cancer classification dataset. Each class is divided into 5 disjoint sets for 5 episodes. The first set, regarded as Domain1 is without any alteration and the next 4 sets for each class undergo various changes to create domain shift. A short description of 4 domain shifts is as follows: Domain 2 (increased stain intensity, simulating, for example, concentration increase of the eosin and/or hematoxylin solutions, each with a different extent of change). Domain 3 (decreased eosin stain intensity, simulating, for example, slides prepared from many years ago with fading stain). Domain 4 (change of hue, simulating, for example, change of reagent manufacturer, scanner or stainer) Domain 5 (change of hue and saturation, simulating, for example, change of reagent manufacturer, scanner or stainer). finally, authors used this augmented dataset for setting class-incremental, domain-incremental, data-incremental, as well as task-incremental scenarios.
@@
Another line of research popularly uses MNIST or similar well-defined datasets instead of collecting real datasets for the evaluation of the CL model. In the medical field, there is a recently released (2021) MNIST-like collection of biomedical images offering dozens of datasets, termed MedMNIST Yang et al. (2023a). Derakhshani et al. (2022) consider 4 MedMNIST datasets including BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST as 4 domains and curate three CL scenarios with it. These datasets are multi-class (8-11 classes in each) disease classification datasets from different imaging modalities and organs. Each dataset was split to have disjoint classes leading to a total of 4 episodes in each dataset. Task and class incremental scenarios were set up for each dataset separately. In the task incremental scenario, the aim was to evaluate only the classes from the current episode whereas in class incremental, the aim was to evaluate all seen classes. Thus the class incremental scenario is difficult than task incremental. They term the datasets generated from different sites/imaging protocols as cross-domain incremental scenario, instead of domain-incremental. Here the evaluation is done across the 4 datasets (BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST), i.e., each dataset is treated as an episode. Further authors presented domain-agnostic and domain-aware settings in this. Domain-agnostic case is more difficult than domain-aware as the domain ID is not explicitly provided in domain-agnostic.
@@ 4 Continual learning strategy
Owing to frequently encountered domain shifts in medical applications, CL strategies aim to tackle performance drop due to domain shifts and thus might become a necessity in clinical applications Pianykh et al. (2020); Lee and Lee (2020). All the available strategies for preventing catastrophic forgetting can be broadly categorized into three main categories: (a) rehearsal-based approaches where a small memory is used to store previous episode data in some form, (b) regularization-based methods aim to control weight update to minimize forgetting previous learning, (c) architectural-based methods mainly aim to keep some network parameters isolated for each episode, and (d) hybrid category which offers various combination of any of these three categories. Figure 6 presents a taxonomy of CL techniques and their adaptation in the medical field by various works. Further, the sub-technique category for each work is also mentioned in the figure. Further, a pie chart showing the ratio of works falling in the above-mentioned categories is shown via Figure 5. It indicates the popularity of using one or more strategies together in hybrid settings owing to better classification or segmentation performances in sequential learning. Now we discuss each of the categories in detail in the following sections.
@@
Hofmanninger et al. (2020a)
Chest CT classification: DIS (3 Epi.)
Experience replay with a dynamic memory size where samples are ranked by computing Gram matrix on last activation maps
Srivastava et al. (2021)
Chest x-ray classification: DIS(3 Epi.)
Latent replay
Perkonigg et al. (2021)
Brain age estimation: DIS (3 Epi.)
Experience replay
Perkonigg et al. (2022)
DIS: three separate experiments for 3 applications: cardiac segmentation (4 Epi.), lung nodule detection (4 Epi.), and brain age estimation (4 Epi.)
Experience replay
Karthik et al. (2022)
DIS on MRI data for brain sclerosis lesions segmentation (8 Epi.)
Random samples from each episode kept for experience replay
Li et al. (2022)
DIS: CMR based Cardiac segmentation across 4 scanners
Generative replay: style-oriented replay
Bera et al. (2023)
TIS (binary segmentation of prostate, spleen, and hippocampus), DIS (binary hippocampus segmentation across 2 datasets), DIS (binary prostate segmentation across 4 datasets)
Experience replay based rehearsal strategy: store max K exemplars (half of the exemplars based on fraction of positive class content inside the image and other half based on gradient variation)
Wang et al. (2023b)
(a) Segmentation for endoscopy: CIS (2-3 Epi.) (b) surgical instrument segmentation (2 Epi.)
Generative replay
@@ 4.1 Rehearsal based
This category of methods aims to approximate and recover old data distributions to augment with the new episode data. Typically there is a memory buffer to store data samples from past episodes which are then used for replaying during the learning of a new episode in order to mitigate catastrophic forgetting of previously learnt knowledge. Here the samples can be original images Rebuffi et al. (2017); Karthik et al. (2022) or deep features Van de Ven and Tolias (2019) or generated pseudo samples Shin et al. (2017) and can be selected via various heuristics and stored in memory buffer.
@@ 4.1.1 Experience replay-based
For natural image classification, Rebuffi et al. (2017) proposed a popular experience-based method called incremental Classifier and Representation Learning (iCaRL) which is also highly explored for various medical applications, e.g., histopathology tumor classification Kaustaban et al. (2022), disease classification Derakhshani et al. (2022), etc. The iCaRL has a fixed memory buffer condition where a subset of the most representative examples is maintained for each class, aiming to carry the most representative information of the class in the learned feature space. The distance between data instances in the latent feature space is used to update the memory buffer. During representation learning, both the stored samples and current task samples are utilized for training. During inference, a nearest-mean-of-exemplars classification strategy is used to assign the label to the class with the most similar prototype. The original iCaRL method requires all data from the new task to be trained together. To address this limitation and enable the new instances from a single task to come at different time steps, Chaudhry et al. (2019) proposed Experience Replay (ER), which uses reservoir sampling to randomly sample a certain number of data instances from a data stream of unknown length, and store them in the memory buffer.
There has been further exploration to select the most representative samples for replay in medical applications. Bera et al. (2023) propose a simple sample selection technique for binary segmentation with a memory bank of $K$ samples, equi-distributed over the classes. Half of the exemplars are selected based on the occupancy of the positive class inside the image, i.e., the higher the content of the positive class, the higher the chance of being selected; rest other half are selected based on their contribution to the learning process, i.e., if the gradient variation is more for a sample then it is hard to learn and hence important. Hofmanninger et al. (2020a) propose to store samples based on their degree of uniqueness inferred using the Gram matrix computed on activations from the last convolution layer of the deep model. They evaluated the framework on chest CT data with synthetically generated domain shifts for classification application. Further, Perkonigg et al. (2021, 2022) propose to combine rehearsal-based continual learning and active learning which indicates important samples that need to be annotated rather than annotating all the samples. They detect domain shifts using a memory buffer for outliers, prompting the labeling of informative samples for model adaptation.
@@ 4.1.2 Latent replay-based
While experience replay requires storing the past samples in raw form, it can cause serious privacy violations in critical medical applications. Therefore a less concerning direction i.e., storing features instead of raw images is also explored. For continual chest x-ray classification application in domain shift conditions, Srivastava et al. (2021) explore leveraging vector-quantization to store and replay hidden representations under memory constraints.
@@ 4.1.3 Generative replay-based
This category of approaches emerged as an alternative to experience replay which highly violates privacy concerns with storing past samples and high memory buffers demands. Here instead of a memory buffer of actual samples, there is a generative model that can generate samples, latent representation, both, or other information related to past tasks. This category of methods is also related to continual learning of generative models where their incremental update is required. Hence, an additional requirement here is the continuous updating of the generative adversarial network. Since actual samples are not stored for replay, this category of approaches is also termed pseudo-rehearsal-based approaches.
To generate the past data, Li et al. (2022) employ a style-oriented replay module which includes a base generative model trained on the first arrived domain and a style bank to record style adjustments for successive domains. Then they incorporate the replayed past data to jointly optimize the model with current data to alleviate catastrophic forgetting. For continual semantic segmentation of MRI data emerging from different institutions, Memmel et al. (2021) train a model by using all the simultaneously available datasets. Their method disentangles content from domain information through adversarial training, resulting in domain-invariant content representations. Other methods follow sequential arrival of a dataset, i.e., one dataset at a time whereas this work requires at least two datasets to start the feature disentanglement learning.
@@ 4.1.4 Memory buffer optimization
Experience replay-based approaches might overfit the stored sub-samples and seem to be bounded by joint training. Alternatively, memory buffer optimization-based approaches following constrained optimization provide solutions that have more scope for both plasticity and stability. The core idea is to perform gradient projection. For example, the Gradient Episodic Memory (GEM) Lopez-Paz and Ranzato (2017) approach corrects the gradient computed on a mini-batch during stochastic gradient descent by utilizing an exemplar set. This prevents changes that could degrade the performance of the network on the exemplar set. Further, the A-GEM approach Chaudhry et al. (2018a) relaxes the projection in one direction computed by randomly selected samples from the memory buffer where both the buffer and sampling size are tunable hyper-parameters. There are some early attempts in this area in the medical realm. For example, for the fundus disease diagnosis problem, Shu et al. (2022) present a gradient regularisation approach for preventing forgetting and a replay-oriented consistency calculation method combined with a subspace weighting strategy to promote model adaptability. The authors employ gradient regularisation in conjunction with a replay-oriented technique. The replay-oriented technique adaptively improves the update for incremental domains without carefully choosing exemplars for replay, whereas gradient regularisation preserves information from previous ones.
@@
Baweja et al. (2018)
MRI based normal brain structures segmentation: TIS (2 Ep.)
EWC
McClure et al. (2018)
Axial and sagittal brain segmentation DIS (4 Ep.)
Distributed Weight Consolidation (DWC)
van Garderen et al. (2019)
MR data based glioma segmentation: DIS (2 epi)
EWC
Özgün et al. (2020)
Brain MRI segmentation: DIS (4 epi)
Learning rate regularization built over memory aware synapses technique
Chen and Tang (2022)
CIS for histopathology-based breast cancer classification (4 Epi)
Regularization: EWC
Zhang et al.; Zhang et al. (2023d; 2021)
DIS for (a) MRI based prostate segmentation, (b) DIS for retinal image based optic cup and disc segmentation
Shape and semantics-based selective regularization to penalize changes of parameters with high joint shape and semantics-based importance
Zhu et al. (2023)
(a) DIS: binary prostate segmentation from T2-weighted MRI scans collected from 6 sites, (b) cross-site and cross-modality 4-class abdominal segmentation between CT and MRI scans, (c) same-site cross-modality muscle segmentation of muscles and bones between MRI and mDixon scans
Knowledge distillation
Liu et al. (2023a)
Incremental brain tumor segmentation: CIS with domain shift (3 epi)
Pseudo label-based knowledge distillation
@@ 4.2 Regularization-based
Rehearsal methods are quite popular due to comparative better performance than other categories; however, the assumption for the availability of past data makes it less suitable for a medical application. In contrast, Regularization-based approaches Kirkpatrick et al. (2017); Schwarz et al. (2018); Zenke et al. (2017) avoid storing examples and mainly add a regularization term in the loss function or regularize the learning rate to penalize model updates that could lead to large deviation from an existing model, thus avoiding forgetting of learned knowledge.

### v3

@@ 3.5 Simulated or Hybrid Settings
It may not be always the case that the model always observes a particular kind of incremental scenario, for example, domain incremental condition. The CL model developed for handling domain incremental conditions may not perform well for a class incremental condition. Thus, there are attempts to evaluate the same model on different CL scenarios; however, a separate evaluation is followed for each kind of incremental CL scenario. In contrast, having a set of episodes with a mix of CL scenarios is close to real-life conditions. Thus it becomes important to also test the performance in hybrid settings. An attempt in direction is made by Sadafi et al. (2023) where other than pure domain incremental and class incremental, a hybrid incremental scenario is also considered. The authors create a sequence of episodes with both class and domain-incremental cases on multi-site WBC classification datasets. They create a long sequence of episodes with a total of 12 episodes. First, they set up a class incremental scenario within the Matek-19 Matek et al. (2019) dataset by adding 3-4 classes in each subsequent episode up to a total of 4 episodes. After this a new domain, i.e., the INT-20 dataset was considered in the same fashion, contributing 4 more episodes. After this, a third domain i.e., a new dataset Acevedo-20 Acevedo et al. (2020) was introduced with 4 episodes in class incremental fashion.
@@
Further, there are some attempts to curate shifts in datasets and various incremental scenarios instead of collecting actual datasets reflecting the situation. For example, Kaustaban et al. (2022) simulated domain shift in H&E stain exhibiting distinct appearance caused by different staining protocols. This dataset is designed to simulate real-world scenarios, where data shifts occur due to differences in scanners, stainers, reagents, and other factors. They considered 9 classes from CRC Kather et al. (2019) which is a H&E stain-based colorectal cancer classification dataset. Each class is divided into 5 disjoint sets for 5 episodes. The first set, regarded as Domain1 is without any alteration and the next 4 sets for each class undergo various changes to create domain shift. A short description of 4 domain shifts is as follows: Domain 2 (increased stain intensity, simulating, for example, concentration increase of the eosin and/or hematoxylin solutions, each with a different extent of change). Domain 3 (decreased eosin stain intensity, simulating, for example, slides prepared from many years ago with fading stain). Domain 4 (change of hue, simulating, for example, change of reagent manufacturer, scanner or stainer) Domain 5 (change of hue and saturation, simulating, for example, change of reagent manufacturer, scanner or stainer). finally, authors used this augmented dataset for setting class-incremental, domain-incremental, data-incremental, as well as task-incremental scenarios.
@@
Another line of research popularly uses MNIST or similar well-defined datasets instead of collecting real datasets for the evaluation of the CL model. In the medical field, there is a recently released (year 2021) MNIST-like collection of biomedical images offering dozens of datasets, termed MedMNIST Yang et al. (2023a). Derakhshani et al. (2022) consider 4 MedMNIST datasets, including BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST as 4 domains and curate 3 CL scenarios with it. These datasets are multi-class (8-11 classes in each) disease classification datasets from different imaging modalities and organs. Each dataset was split to have disjoint classes leading to a total of 4 episodes in each dataset. Task and class incremental scenarios were set up for each dataset separately. In the task incremental scenario, the aim was to evaluate only the classes from the current episode whereas in class incremental, the aim was to evaluate all seen classes. Thus the class incremental scenario is difficult than task incremental. They term the datasets generated from different sites/imaging protocols as cross-domain incremental scenario, instead of domain-incremental. Here the evaluation is done across the 4 datasets (BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST), i.e., each dataset is treated as an episode. Further authors presented domain-agnostic and domain-aware settings in this. Domain-agnostic case is more difficult than domain-aware as the domain ID is not explicitly provided in domain-agnostic.
@@ 4 Continual learning strategy
Owing to frequently encountered domain shifts in medical applications, CL strategies aim to tackle performance drop due to domain shifts and thus might become a necessity in clinical applications Pianykh et al. (2020); Lee and Lee (2020). All the available strategies for preventing catastrophic forgetting can be broadly categorized into three main categories: (a) rehearsal-based approaches where a small memory is used to store previous episode data in some form, (b) regularization-based methods where the aim is to control weight update to minimize forgetting the previous learning, (c) architectural-based methods which mainly aim to keep some network parameters isolated for each episode, and (d) hybrid category which offers various combination of any of these three categories. Figure 6 presents a taxonomy of CL techniques and their adaptation in the medical field by various works. Further, the sub-technique category for each work is also mentioned in the figure. Additionally, a pie chart showing the ratio of works falling in the above-mentioned categories is shown via Figure 5. It indicates the popularity of using one or more strategies together in hybrid settings owing to better classification or segmentation performances in sequential learning. Now, we discuss each category in detail in the following sections.
@@
Hofmanninger et al. (2020a) Chest CT classification: DIS (3 Epi.) Experience replay with a dynamic memory size where samples are ranked by computing Gram matrix on last activation maps
Srivastava et al. (2021) Chest x-ray classification: DIS(3 Epi.) Latent replay
Perkonigg et al. (2021) Brain age estimation: DIS (3 Epi.) Experience replay
Perkonigg et al. (2022) DIS: three separate experiments for 3 applications: cardiac segmentation (4 Epi.), lung nodule detection (4 Epi.), and brain age estimation (4 Epi.) Experience replay
Karthik et al. (2022) DIS on MRI data for brain sclerosis lesions segmentation (8 Epi.) Random samples from each episode kept for experience replay
Li et al. (2022) DIS: CMR based Cardiac segmentation across 4 scanners Generative replay: style-oriented replay
Bera et al. (2023) TIS (binary segmentation of prostate, spleen, and hippocampus), DIS (binary hippocampus segmentation across 2 datasets), DIS (binary prostate segmentation across 4 datasets) Experience replay based rehearsal strategy: store max K exemplars (half of the exemplars based on fraction of positive class content inside the image and other half based on gradient variation)
Wang et al. (2023b) (a) Segmentation for endoscopy: CIS (2-3 Epi.) (b) surgical instrument segmentation (2 Epi.) Generative replay
@@ 4.1 Rehearsal based
This category of methods aims to approximate and recover old data distributions to augment with the new task data. Typically, there is a memory buffer to store data samples from past tasks, which are then used for replaying during the learning of a new task in order to retain previously learned knowledge and hence mitigate catastrophic forgetting. Here, the samples can be original images Rebuffi et al. (2017); Karthik et al. (2022) or deep features Van de Ven and Tolias (2019) or generated pseudo samples Shin et al. (2017) and can be selected via various heuristics and stored in memory buffer.
@@ 4.1.1 Experience replay-based
For natural image classification, Rebuffi et al. (2017) proposed a popular experience-based method called incremental Classifier and Representation Learning (iCaRL) which is also highly explored for various medical applications, e.g., histopathology tumor classification Kaustaban et al. (2022), disease classification Derakhshani et al. (2022), etc. The iCaRL strategy has a fixed memory buffer condition where a subset of the most representative examples is maintained for each class, aiming to carry the most representative information of the class in the learned feature space. The distance between data instances in the latent feature space is used to update the memory buffer. During representation learning, both the stored samples and current task samples are utilized for training. During inference, a nearest-mean-of-exemplars classification strategy is used to assign the label to the class with the most similar prototype. The original iCaRL method requires all data from the new task to be trained together. To address this limitation and enable the new instances from a single task to come at different time steps, Chaudhry et al. (2019) proposed Experience Replay (ER), which uses reservoir sampling to randomly sample a certain number of data instances from a data stream of unknown length, and store them in the memory buffer.
Further exploration has been conducted to select the most representative samples for replay in medical applications. Bera et al. (2023) propose a simple sample selection technique for binary segmentation with a memory bank of $K$ samples, equi-distributed over the classes. Half of the exemplars are selected based on the occupancy of the positive class inside the image, i.e., the higher the content of the positive class, the higher the chance of being selected; the other half are selected based on their contribution to the learning process, i.e., if the gradient variation is more for a sample, then it is hard to learn and hence important. Hofmanninger et al. (2020a) propose to store samples based on their degree of uniqueness inferred using the Gram matrix computed on activations from the last convolution layer of the deep model. They evaluated the framework on chest CT data with synthetically generated domain shifts for classification application. Further, Perkonigg et al. (2021, 2022) propose to combine rehearsal-based CL and active learning, where the model indicates important samples that need to be annotated rather than annotating all the samples. They detect domain shifts using a memory buffer for outliers, prompting the labeling of informative samples for model adaptation.
@@ 4.1.2 Latent replay-based
While experience replay requires storing the past samples in raw form, it can cause serious privacy violations in critical medical applications. Therefore, a less concerning direction, i.e., storing features instead of raw images, is also explored. For continual chest x-ray classification application in domain shift conditions, Srivastava et al. (2021) explore leveraging vector-quantization to store and replay hidden representations under memory constraints.
@@ 4.1.3 Generative replay-based
This category of approaches emerged as an alternative to experience replay, which highly violates privacy concerns with storing past samples and high memory buffer demands. Here, instead of a memory buffer of actual samples, there is a generative model that can generate samples, latent representation, both, or other information related to past tasks. This category of methods is also related to incremental learning of generative models where their incremental update is required. Hence, an additional requirement here is the continuous updating of the generative adversarial network. Since actual samples are not stored for replay, this category of approaches is also termed pseudo-rehearsal-based approaches.
To generate the past data, Li et al. (2022) employ a style-oriented replay module, which includes a base generative model trained on the first arrived domain and a style bank to record style adjustments for successive domains. Then, they incorporate the replayed past data to jointly optimize the model with current data to alleviate catastrophic forgetting. For continual semantic segmentation of MRI data emerging from different institutions, Memmel et al. (2021) train a model by using all the simultaneously available datasets. Their method disentangles content from domain information through adversarial training, resulting in domain-invariant content representations. Other methods follow the sequential arrival of a dataset, i.e., one dataset at a time, whereas this work requires at least two datasets to start the feature disentanglement learning.
@@ 4.1.4 Memory buffer optimization
Experience replay-based approaches might overfit the stored sub-samples and seem to be bounded by joint training. Alternatively, memory buffer optimization-based approaches following constrained optimization provide solutions that have more scope for both plasticity and stability. The core idea is to perform gradient projection. For example, the Gradient Episodic Memory (GEM) Lopez-Paz and Ranzato (2017) approach corrects the gradient computed on a mini-batch during stochastic gradient descent by utilizing an exemplar set. This prevents changes that could degrade the performance of the network on the exemplar set. Further, the A-GEM approach Chaudhry et al. (2018a) relaxes the projection in one direction computed by randomly selected samples from the memory buffer where both the buffer and sampling size are tunable hyper-parameters. There are some early attempts in this area in the medical realm. For example, for the fundus disease diagnosis problem, Shu et al. (2022) presents a gradient regularisation approach for preventing forgetting and a replay-oriented consistency calculation method combined with a subspace weighting strategy to promote model adaptability. The authors employ gradient regularisation in conjunction with a replay-oriented technique. The replay-oriented technique adaptively improves the update for incremental domains without carefully choosing exemplars for replay, whereas gradient regularisation preserves information from previous ones.
@@
Baweja et al. (2018) MRI based normal brain structures segmentation: TIS (2 Ep.) EWC
McClure et al. (2018) Axial and sagittal brain segmentation DIS (4 Ep.) Distributed Weight Consolidation (DWC)
van Garderen et al. (2019) MR data based glioma segmentation: DIS (2 epi) EWC
Özgün et al. (2020) Brain MRI segmentation: DIS (4 epi) Learning rate regularization built over memory aware synapses technique
Chen and Tang (2022) CIS for histopathology-based breast cancer classification (4 Epi) Regularization: EWC
Zhang et al.; Zhang et al. (2023d; 2021) DIS for (a) MRI based prostate segmentation, (b) DIS for retinal image based optic cup and disc segmentation Shape and semantics-based selective regularization to penalize changes of parameters with high joint shape and semantics-based importance
Zhu et al. (2023) (a) DIS: binary prostate segmentation from T2-weighted MRI scans collected from 6 sites, (b) cross-site and cross-modality 4-class abdominal segmentation between CT and MRI scans, (c) same-site cross-modality muscle segmentation of muscles and bones between MRI and mDixon scans Knowledge distillation
Liu et al. (2023a) Incremental brain tumor segmentation: CIS with domain shift (3 epi) Pseudo label-based knowledge distillation
@@ 4.2 Regularization-based
Rehearsal methods are quite popular due to their comparative better performance than other categories; however, the assumption of the availability of past data makes them less suitable for medical applications. In contrast, Regularization-based approaches Kirkpatrick et al. (2017); Schwarz et al. (2018); Zenke et al. (2017) avoid storing examples and mainly add a regularization term in the loss function or regularize the learning rate to penalize model updates that could lead to large deviation from an existing model, thus avoiding forgetting of learned knowledge.